# Laplace comparison on Kähler Ricci flow and convergence

### v1

@@ 1. Introduction
Given initial Kähler metric $g_{i\bar{j}}(0)$, it is proved in [Ca] that (1.1) has a solution for all time $t$. Recently, many results concerning long time and uniform behavior of (1.1) have appeared. For example, when the curvature operator or the bisectional curvature is nonnegative, it is known that solutions to (1.1) stays smooth when time goes to infinity (see [CCZ] for examples). In the general case, Perelman (cf [ST]) proved that the scalar curvature $R$ is uniformly bounded, and the Ricci potential $u(\cdot,t)$ is uniformly bounded in $C^{1}$ norm, with respect to $g(t)$. When the complex dimension $m=2$, let $(M,g(t))$ be a solution to (1.1), it is proved in ([CW]) that the flow sequentially converges to an orbifold. When $m=3$, the authors of [TZz] proved sequential convergence except on a small singular set of co-dimension 4. For the general case, in the paper [TZq1], it was proved that a Gromov-Hausdorff limit of the flow is a metric space with volume doubling and $L^{2}$ Poincaré inequality and $L^{1}$ isoperimetric inequality. In [LTZ], it is proven that the curvature tensor for some KRF can diverge to infinity as time goes to $\infty$. The convergence issue of Kähler Ricci flows have attracted much attention recently. See for example [MS], [PS1], [PSSW1], [PSSW2], [Se], [SoT], [Sz], [SW], [To], [TZhu1], [TZhu2], [TZZZ], [Zz1], [Zz2] and [Zhu] and references therein. These activities are centered around the:
@@
As mentioned above, this conjecture was confirmed in complex dimensions 2 and 3 in [CW] and [TZz] respectively. In the higher dimension cases, several years ago, two substantial papers [CW4] and [Ba] have appeared where solutions to the Hamilton-Tian conjecture are presented. However, there are some arguments in these papers which seem to be challenging for us, more specifically, we thought that the key assumptions in the induction process in these papers need further clarification for its validity. We refer the readers to the Appendix for details. In a more recent paper [WZ2], another interesting path to this conjecture is presented, based on two ingredients. The first one is the lower bound of the Bergman functions for (1.1) proven in [Zhkw], based on [LS]. The other is the Kodaira embedding of $M$ into $CP^{N}$ by sections of holomorphic line bundles. It appears that the co-dimension 4 property of the singular set still needs a little extra work except for the Kähler Einstein case.
@@ 6. There exists a uniform constant $S_{2}$ so that the following $L^{2}$ Sobolev inequality holds:
Properties 1-3 and 4 are proven by Perelman (c.f. [ST], [P]); Property 5 can be found in [Z11] and also [CW2]; Property 6 was first proven in [Z07] (see also [Ye], [Z10] ). Properties 7 and 8 are in [TZq1], and Property 9 is in [TZq2]. Property 10 is in [TZz].
@@
Here $\rho_{0}\geq 0$ is a constant depending only on the main parameters of the initial metric. In the above, we have used the fact that the diameters of $(M,g(t))$ are uniformly bounded due to Perelman c.f. [ST]. In another word, the quadratic form on the left hand side is bounded from below. Let us choose $\rho_{0}$ to be the optimal one such that (2.22) holds, i.e.
@@ 3.2. $\epsilon$ splitting map, metric cone property, codimension 4 singular set
By scaling, it is enough to consider the scaled metric $\hat{g}=r^{-2}\tilde{g}$ and set $r=1$. For simplicity of presentation, we will drop the hat from $\hat{g}$ in the rest of the proof. So $g$ actually means $\hat{g}$. Also, if no confusions arise will will omit the reference to the metric in all geometric quantities, unless stated otherwise. In addition all the constants depend only on the basic parameters of $g(0)$
@@
To extend this result to $(M,\tilde{g}(t_{i}))$, there are apparently two obstacles. The first is the that $(M,\tilde{g}(t_{i}))$ is not Kähler. The second is that we do not know the Ricci curvature is uniformly bounded. However since the conformal factors satisfy $\|h_{i}\|_{C^{1/2}}\leq C$, the tangent cones of $(X,d)$ are Kähler since they are the same as those of the limit space of the Kähler manifolds $(M,g(t_{i}))$. Therefore we only need to prove the conclusion of Theorem 1.22 in [CCT] which is the corresponding real version of Theorem 1.23. By checking the proof of this theorem, after establishing the $\epsilon$ splitting and metric cone properties and Lemma 2.18, there are only two places where the boundedness of the Ricci curvature is needed. One is the proof of Theorem 3.7 which uses (3.6) that involves the Ricci curvature, the other is in the proof of Theorem 7.2 where the existence of $C^{1,\alpha}$ harmonic coordinates in the sense of [An] is needed. Fortunately, the a priori bounds for $Ric_{\tilde{g}}$ in the suitable Morrey and Kato class in Proposition 3.1 is enough to carry out the proof.
@@
4. Appendix
In this section, we aim to describe some possible issues in two previous publications.
1. lower bound of the Bergman kernel goes to -$\infty$ when the metrics blow up in [CW4].
Let $(M,g(t))$ be as in (1.1). One calls $LM=\{(M,g(t),J,L,h(t)),t\in[a,b]\subset\mathbf{R}^{1}\}$ a polarized Kähler Ricci flow if $J$ is the complex structure, $L$ is a pluri anti-canonical line bundle over $M$, $h(t)$ is a family of smooth metrics on $L$ whose curvature is $\omega(t)$, the metric form compatible with $g(t)$ and the complex structure $J$.
For a large positive integer $k$, the Bergman function $b^{(k)}$ is defined as
$b^{(k)}(x,t)=\log\sum^{N_{k}}_{i=0}\|S^{(k)}_{i}\|^{2}_{h(t)}(x,t).$ (4.1)
Here $N_{k}=dim_{\mathbb{C}}H^{0}(M,L^{k})-1$, $\{S^{(k)}_{i}\}$ are orthonormal basis of $H^{0}(M,L^{k})$ under the metrics $\omega(t)$ and $h(t)$.
In the definition that the polarized canonical radius (pcr) being bounded from below by a positive constant, say $1$, on p36 [CW4], one of the assumptions is a definite lower bound of the Bergman kernel
$\sup_{1\leq j\leq 2k_{0}}b^{(j)}(x,t)\geq-2k_{0}.$ (4.2)
Note the sup is on $j$ not space variable $x$ nor $t$. For the KRF (1.1), one does not know if the pcr is bounded from below by a definite constant, say $1/j_{0}$. In order to prove it, the authors use a blow up argument on p81. In (3.127), line 2, it is stated that for the Ricci flow scaled by $4kj^{2}_{0}$, the pcr is bounded from below by $1/2$. Here $k$ will go to infinity and $j_{0}$ is fixed.
However, a direct computation by scaling shows that the infimum of the Bergman function of the scaled manifold goes to -$\infty$ as $k\to\infty$. So the pcr can not be bounded from below by a positive constant which is independent of $k$ or $k_{0}$. Indeed, let $\tilde{g}=4kj^{2}_{0}g$ be the scaled metric and $\tilde{b}^{(j)}$, $j=1,...,k_{0}$ be the scaled Bergman function. Then, by definition,
$\tilde{b}^{(j)}(x,t)=\log\sum^{N_{j}}_{i=0}\|\tilde{S}^{(j)}_{i}\|^{2}_{\tilde{h}(t)}(x,t).$ (4.3)
Here $\{\tilde{S}^{(j)}_{i}\}$ are orthonormal basis of $H^{0}(M,L^{j})$ under the metrics $\tilde{\omega}(t)=\tilde{g}$ and $\tilde{h}(t)$. Therefore
$\int_{M}\sum^{N_{j}}_{i=0}\|\tilde{S}^{(j)}_{i}\|^{2}_{\tilde{h}(t)}(x,t)d\tilde{g}\leq CN_{j}.$ (4.4)
Here $C$ is a constant independent of $k$ or $j$. Note $j\leq k_{0}$ which is fixed. Therefore
$\inf_{M}\sum^{N_{j}}_{i=0}\|\tilde{S}^{(j)}_{i}\|^{2}_{\tilde{h}(t)}(x,t)\leq C\frac{N_{j}}{|M|_{\tilde{g}}}\to 0,\qquad k\to\infty.$ (4.5)
Here we just used the relation on volumes of the manifold $M$ under $g$ and $\tilde{g}$
$|M|_{\tilde{g}}=(4kj^{2}_{0})^{n/2}|M|_{g}\to\infty.$
Hence, by (4.3),
$\inf_{M}\tilde{b}^{(j)}(x,t)\to-\infty.$ (4.6)
In fact, by Chebyshev inequality, it is easy to see that the set where the Bergman function is bounded from below by a fixed constant has a finite measure. So in most places, the Bergman function is not bounded from below when the metric is blown up. The same also hold for any finite sum of $|b^{(j)}|$ as long as the number of terms is bounded.
2. The conclusion of Lemma 6.3 in [Ba] seems non-uniform.
On p821 of the cited paper, the following key estimate for curvatures under a uniform scale $\bar{r}$ was given.
Lemma 6.3. For any $A<\infty$ and $0<p<4$, there are constants $E^{\prime}=E^{\prime}(A)<\infty$ and $\bar{r}=\bar{r}(A)$ such that the following holds.
Let $(M,(g(t))_{t\in[0,2]})$ be a Ricci flow on a compact $n$ dimensional and orientable manifold $M$ with the property that
(i) The ”initial” $\nu$ entropy at scale $4$ satisfies: $\nu[g(-2),4]\geq-A$;
(ii) $|R|\leq 1$ on $M\times[-2,0]$.
Then for any $(x,t)\in M\times[-1,0]$ and $0<r\leq\bar{r}$ and $0<s<1$, we have
$\left|\{r_{Rm}(\cdot,0)<sr\}\cap B(x,r,g(0))\right|_{g(0)}\leq E^{\prime}s^{3.1}r^{n}.$ (4.7)
Here
$r_{Rm}(x,t)=\sup\{\rho>0\,|\,\rho^{2}\|Rm(\cdot,t)\|_{L^{\infty}(B(x,\rho,g(t)))}<1\}$
is the curvature radius.
We mention that (4.7) holds automatically under a non-uniform scale just by smoothness of the Ricci flow. For example, if $\bar{r}$ is sufficiently small, then the left hand side of (4.7) is already 0. So the uniformity of $\bar{r}$ is crucial.
The proof of the lemma is based on Lemma 6.2 on p820, which is also listed below for convenience. By checking the proof of Lemma 6.3, it seems to us that the radius $\bar{r}$ depends on the Ricci flow itself instead of only on the entropy lower bound $A$ as stated.
The reason is that in the first line of the proof, it is written that ”Fix $A$ and choose $E^{\prime}:=\mathbf{E}_{3.1}(A)<\infty$ according to Lemma 6.2. Next choose $0<\bar{r}=\bar{r}(A,3,1,E^{\prime})<1$ according to Lemma 6.2.” Then on the next page (p822), 2nd paragraph in italic, it is stated that if (4.7) fails in a parabolic cube of size $r$, then it fails in a parabolic cube of size $r/10$. This is due to Lemma 6.2 listed below, stated in the contrapositive way.
However, in condition (iii) of Lemma 6.2, the constant $E^{\prime}$ can be arbitrary. It may not be the $\mathbf{E}_{3.1}(A)$ that was named $E^{\prime}$ to begin with. Also the radius $\bar{r}$ in Lemma 6.2 depend on the $E^{\prime}$ in the upper bound in condition (iii) of that lemma. The issue is that one does not know how large it is a priori. In general one can not choose it as a given fixed number such as $\mathbf{E}_{3.1}(A)$ stated in the proof. Therefore it is not certain that the iteration in the next paragraph on p822 can be continued. For example, if the contrapositive $E^{\prime}$ in (iii) of Lemma 6.2 is smaller than $\mathbf{E}_{3.1}(A)$ in one of the steps, then the iteration can not continue.
Lemma 6.2 ([Ba]). For any $A<\infty$ and $0<p<4$, there is a constant $E_{p}(A)<\infty$ such that for any $E^{\prime}<\infty$, there is a constant $0<\bar{r}=\bar{r}(A,p,E^{\prime})<1$ such that the following holds.
Let $(M,(g(t))_{t\in[0,2]})$ be a Ricci flow on a compact $n$ dimensional orientable manifold $M$ and $0<r_{0}\leq\bar{r}$ a scale with the property that
(i) The ”initial” $\nu$ entropy at scale $4$ satisfies: $\nu[g(-2),4]\geq-A$.
(ii) $|R|\leq 1$ on $M\times[-2,0]$;
(iii) for all $(x,t)\in M\times[-1,0]$ and $0<r<r_{0}$ and $0<s<1$, we have
$\left|\{r_{Rm}(\cdot,t)<sr\}\cap B(x,r,g(t))\right|_{g(t)}\leq E^{\prime}s^{3.1}r^{n}.$
Then for any $x\in M$ and $0<r\leq 10r_{0}$ and $0<s<1$, we have
$\left|\{r_{Rm}(\cdot,0)<sr\}\cap B(x,r,g(0))\right|_{g(0)}\leq E_{p}s^{p}r^{n}.$ (4.8)
To recap the above discussion, one assumes (i) and (ii) hold and $p=3.1$ as stated. Then the contrapositive statement of Lemma 6.2 is: if (4.8) fails for some $x,r\leq 10r_{0},s$, then (iii) fails some some $E^{\prime}>0$, $x,r<r_{0},s$. But this $E^{\prime}$ may be smaller than $E_{3.1}$.
@@ Acknowledgements
Q. S. Z. gratefully acknowledges the support of Simons’ Foundation grant 710364. Thanks also go to Professors R. Bamler, X.X. Chen, B. Wang and Guofang Wei for helpful comments.

### v2

@@ 1. Introduction
Given initial Kähler metric $g_{i\bar{j}}(0)$, it is proved in [Ca] that (1.1) has a solution for all time $t$. Recently, many results concerning long time and uniform behavior of (1.1) have appeared. For example, when the curvature operator or the bisectional curvature is nonnegative, it is known that solutions to (1.1) stays smooth when time goes to infinity (see [CCZ] for examples). In the general case, Perelman (cf [ST]) proved that the scalar curvature $R$ is uniformly bounded, and the Ricci potential $u(\cdot,t)$ is uniformly bounded in $C^{1}$ norm, with respect to $g(t)$. When the complex dimension $m=2$, let $(M,g(t))$ be a solution to (1.1), it is proved in ([CW]) that the flow sequentially converges to an orbifold. When $m=3$, the authors of [TZz] proved sequential convergence except on a small singular set of co-dimension 4. For the general case, in the paper [TZq1], it was proved that a Gromov-Hausdorff limit of the flow is a metric space with volume doubling and $L^{2}$ Poincaré inequality and $L^{1}$ isoperimetric inequality. In [LTZ], it is proven that the curvature tensor for some KRF can diverge to infinity as time goes to $\infty$. The convergence issue of Kähler Ricci flows have attracted much attention recently. See for example [MS], [PS1], [PSSW1], [PSSW2], [Se], [SoT], [CZ], [Sz], [SW], [To], [TZhu1], [TZhu2], [TZZZ], [Zz1], [Zz2] and [Zhu] and references therein. Much of these activities are centered around the:
@@
As mentioned above, this conjecture was confirmed in complex dimensions 2 and 3 in [CW] and [TZz] respectively. In the higher dimension cases, several years ago, two substantial papers [CW4] and [Ba] have appeared where solutions to the Hamilton-Tian conjecture are presented. See also Section 4. In a more recent paper [WZ2], another interesting path to this conjecture is presented, based on two ingredients. The first one is the lower bound of the Bergman functions for (1.1) proven in [Zhkw], based on [LS]. The other is the Kodaira embedding of $M$ into $CP^{N}$ by sections of holomorphic line bundles. It appears that the co-dimension 4 property of the singular set still needs a little extra work except for the Kähler Einstein case.
@@ 6. There exists a uniform constant $S_{2}$ so that the following $L^{2}$ Sobolev inequality holds:
Properties 1-3 and 4 are proven by Perelman (c.f. [ST], [P]); Property 5 can be found in [Z11] and also [CW2]; Property 6 was first proven in [Z07] (see also [Ye], [Z10] ). Properties 7 and 8 are in [TZq1], proven as a consequence of properties 3, 4, 5, 6, and Property 9 is in [TZq2]. Property 10 is in [TZz].
@@
Note that one can assume that $V$ is smooth without losing generality since one can first choose the function $V_{1}=|Ric^{-}(x)|+\epsilon$ and smooth out $V_{1}$ to a smooth function $V\geq|Ric^{-}(x)|$.
@@
Here $\rho_{0}\geq 0$ is a constant depending only on the main parameters of the initial metric. See also (2.29) below. In the above, we have used the fact that the diameters of $(M,g(t))$ are uniformly bounded due to Perelman c.f. [ST]. In another word, the quadratic form on the left hand side is bounded from below. Let us choose $\rho_{0}$ to be the optimal one such that (2.22) holds, i.e.
@@ 3.2. $\epsilon$ splitting map, metric cone property, codimension 4 singular set
By scaling, it is enough to consider the scaled metric $\hat{g}=r^{-2}\tilde{g}$ and set $r=1$. For simplicity of presentation, we will drop the hat from $\hat{g}$ in the rest of the proof. So $g$ actually means $\hat{g}$. Also, if no confusions arise will omit the reference to the metric in all geometric quantities, unless stated otherwise. In addition all the constants depend only on the basic parameters of $g(0)$
@@
To extend this result to $(M,\tilde{g}(t_{i}))$, there are apparently two obstacles. The first is that $(M,\tilde{g}(t_{i}))$ is not Kähler. The second is that we do not know the Ricci curvature is uniformly bounded. However since the conformal factors satisfy $\|h_{i}\|_{C^{1/2}}\leq C$, the tangent cones of $(X,d)$ are Kähler since they are the same as those of the limit space of the Kähler manifolds $(M,g(t_{i}))$. Therefore we only need to prove the conclusion of Theorem 1.22 in [CCT] which is the corresponding real version of Theorem 1.23. By checking the proof of this theorem, after establishing the $\epsilon$ splitting and metric cone properties and Lemma 2.18, there are only two places where the boundedness of the Ricci curvature is needed. One is the proof of Theorem 3.7 which uses (3.6) that involves the Ricci curvature, the other is in the proof of Theorem 7.2 where the existence of $C^{1,\alpha}$ harmonic coordinates in the sense of [An] is needed. Fortunately, the a priori bounds for $Ric_{\tilde{g}}$ in the suitable Morrey and Kato class in Proposition 3.1 is enough to carry out the proof.
@@
4. Remark
In the first version of the paper, there is an appendix in which we mentioned possible issues in two previous publications.
One is the lower bound of the Bergman kernel when the metrics blow up in [CW4]. Afterwards the authors of [CW4] posted a note on arXiv disputing our claim. We have responded to them in more details about our concerns.
The other one is a possible issue in going from Lemma 6.2 to Lemma 6.3 in [Ba]. Recently the author of [Ba] informed us that the proof of Lemma 6.3 works and explained his arguments. We agree with it and thank professor R. Bamler for his patient explanation.
After further consultation and discussion, the appendix in version 1 of the paper is replaced by a supplement with the link ”https://sites.google.com/view/qiszhang-ucr249/home”, click ” Supplement.arxiv.2509.14820”
@@ Acknowledgements
Q. S. Z. gratefully acknowledges the support of Simons’ Foundation grant 710364. Thanks also go to Professors R. Bamler, X.X. Chen, B. Wang and Guofang Wei for helpful discussion over the years.